If an energy storage system is capable of providing 10 kW at 80 % efficiency, the total energy capacity required to provide 8 hours of operation is

Answer: 100kWh
Answer #1

10 KW or 10,000 joules / sec

in 8 hour total energy

8 hour = 28800 sec

= 10,000 * 28800

= 2.88*10^8 joules

this is 80% energy

efficiency = 80%

so, energy =

= (2.88*10^8 / 80 )*100

3.6*10^8

in Kwh

1 Kwh = 2.8*10^-7

3.6*10^8*2.8*10^-7

= 100.8 KWh

any query please comment

if satisfied please rate it thanks